OpenStudy (anonymous):

solve algebraically 2sin^2 theta-4sin theta=cos^2 theta-2

OpenStudy (loser66):

replace cos^2 = 1-sin^2 let everything in LHS solve quadratic polynomial with sin theta is variable

OpenStudy (loser66):

hey, I guide only, I don't work for you. You do it,

OpenStudy (anonymous):

what you think im doing now

OpenStudy (anonymous):

THANKS REALLY HELPED ALOT

OpenStudy (loser66):

glad to hear that THANKS REALLY HELPED ALOT, LOL, LOL LOL.... , actually what does it mean? sarcastic? why do you shout at me with your appreciation? heheheha

OpenStudy (anonymous):

I SHOUT WITH APPRECIATION BECAUSE I FINALLY UNDERSTAND IT!!!

OpenStudy (loser66):

hahahahaa......

OpenStudy (anonymous):

trig is hard so if i understand one thing im excited

OpenStudy (loser66):

sometimes, trig is trick, too

OpenStudy (anonymous):

yeah sometimes i just want to burn my trig notebook

OpenStudy (kenljw):

2sin^2 theta-4sin theta=cos^2 theta-2 sin^2(x) + sin^2(x) - cos^2(x) +2- 4sin(x) = 0 sin^2(x) + cos^2(x) + sin^2(x) + cos^2(x) - 3cos^2(x) + 2 - 4sin(x) =0 sin^2(x) + cos^2(x) =1 -3[cos^2(x) + (4/3)sin(x) - 4/3] -3[1 - sin^2(x) +(4/3)sin(x) - 4/3] = 3[sin^2(x) - (4/3)sin(x) + 1/3] = 0 set y = sin(x) solve quadratic y^2 - (4/3)y -1/3 = (y - 1/3)(y -1) = 0 sin(x) = y = 1, 1/3 therefore x = pi/2 = 90 deg, 19.47 deg Did anyone else get these answers?